TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Dan Dubrick
date: 2003-07-15 00:48:00
subject: 7\10 Pt 1 ISS On-Orbit Status 10-07-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

13 July 2003

ISS On-Orbit Status 10 Jul 2003

Part 1 of 3

All ISS systems continue to function nominally, except as noted
previously or below.

FE/SO Ed Lu completed the standard general MedOps periodic fitness
evaluation (PFE), checking up on his blood pressure and
electrocardiogram during programmed exercise on the CEVIS in the
Lab.  Readings were taken with the BP/ECG (blood
pressure/electrocardiograph), and his crewmate, CDR Yuri Malenchenko,
assisted as CMO (crew medical officer).  [BP/ECG provides automated
noninvasive systolic and diastolic blood pressure measurements while
also monitoring and displaying accurate heart rates on a continual
basis at rest and during exercise.]

Malenchenko started his work day with another data take of the
Russian MO-21 "Ecosfera" air sampler and incubation equipment,
collecting air samples for atmospheric microbial analysis to monitor
sanitary-hygiene status.  [MO-21 determines microbial contamination
of the ISS atmosphere, specifically the total bacterial and fungal
microflora counts and microflora composition according to morphologic
criteria of microorganism colonies.] 

Ed Lu recharged both batteries of the MedOps defibrillator equipment
to full capacity.  [Each battery was left in for 3.5 hrs. and its
open-circuit voltage tested with the volt/amp scopemeter, then
removed and stowed again. Nominally, the Defib has a battery
installed at all times, but with this particular unit the PDIM (power
data interface module) would overcharge the batteries.]

Yuri Malenchenko concluded his modification work on the Service
Module (SM) condensate separation & pumping units (BRPK).  [After
yesterday's replacement of BRPK-2, today the BRPK units were to be
upgraded with new pipe conduits without filters and with higher trip
pressures for the safety valves (also new).  Purpose of this upgrade
is to prevent unwanted fluid flow through the bypass line during
condensate transfer cycles.] 

Malenchenko then continued the latest maintenance cycle on the
Russian segment (RS) ventilation system, today cleaning the
protective mesh screens of the ventilation fans (TsV) and the
removable screens of the three gas-liquid heat exchangers (GZhT) in
the FGB (funktsionalnyi-grusovoi blok) air ventilation system,
disposing of the old filters.  [Last time done: 6/6].

Yuri also attended to his regular daily inspection of the BIO-5
Rasteniya-2 ("Plants-2") experiment which studies growth and
development of plants (peas) under spaceflight conditions in the
Lada-2 greenhouse.

Ed Lu conducted another weekly inventory audit of the available CWCs
(contingency water containers) and their contents, to keep track of
onboard water supplies.  An updated CWC "cue card" was uplinked to
support this weekly activity with an updated list of residual water
resources remaining in the partially-filled CWCs.  [Last time done:
7/2].

The Science Officer performed the routine maintenance of SM life
support systems (SOZh) and also prepared the daily inventory
management system (IMS) "delta" file for updating the IMS database.

The crew conducted the regular bi-monthly teleconference with the JSC
Astronaut Office (Kent Rominger).

Ed Lu had the second day for his recording of video scenes of a
two-day onboard tour for the "Live on ISS" PAO TV program, started
yesterday.  The footage will be edited by him on 7/11 (tomorrow) and
downlinked on 7/17 (Thursday next week).  The downlink, with live
crew narration, will be aired on NASA TV.  [Scenes of interest
include an end-to-end station "fly-through", starting from the Lab
and ending at the SM, highlights of the Lab interior, the MSG
(microgravity science glovebox), EarthKAM and TVIS treadmill, the
Airlock and its spacesuits, if visible, the FGB and its stores, the
Soyuz TMA-1 CRV, views of outside trusses, etc.] 

At 12:25pm EDT, the two station residents were interviewed live on TV
by Voice of America Television and Radio while floating in the
"Destiny" Lab. The 18-min. event touched on recent activities aboard
ISS, operational impressions by the crew, and return-to-flight
questions. 

Both crewmembers conducted their regular daily physical exercise
program of 2.5 hrs.

Ed Lu's attempts yesterday at recovering the two SSC laptops #6071 &
#6039 were not successful.  After running for a while, first with,
then without battery, both PCs locked up again.  More troubleshooting
by ground and crew ahead.

More success was had by the SO with the restoration of the SAMS ICU
(space acceleration measurement system/interface control unit) laptop
hard drive. The problem was resolved and the unit is working
nominally. 

Malenchenko's work on repairing the Russian laptop #1 was also
successful. 

The BMP harmful impurities filtration system in the SM is back in
operation after the recent IFM.

EarthKAM operations are continuing on board, under remote commanding.

Moscow reported that the installation of the new 11P-delivered
satellite navigation antenna system (ASN) in the SM by Yuri on 7/8
was successfully completed, as were two subsequent tests, conducted
remotely from TsUP. [Telemetry analysis is continuing at present.
Another set of tests is to begin on 7/14, to include several
different operational modes and software alternatives in both the
primary and backup system, to run for about two weeks.  If
successful, the system will then be declared operational.  Using
GLONASS satellites (the Russian equivalent of GPS), the new
capability will allow state vector (SV) updates without using the
ground (which up to now has to uplink daily SV updates) or requiring
SV transfers from the USOS from time to time.  The ASN equipment was
originally installed in the SM but was found faulty and had to be
returned to the ground.  After repair it was shipped again to the
station and re-installed.] 

The degraded condition of the RS power supply situation continues and
is being carefully monitored.  The replacement of the failed SM
battery #4 is planned in the near future.

Tomorrow at 2:00pm EDT, the station will maneuver to YVV/"barbecue"
attitude (vertical y-axis in velocity vector, i.e., facing forward)
for six orbits, to allow gathering temperature and power data for the
scheduled mid-Beta range analysis test.  Afterwards, flight attitude
will return to XPOP. [Projected power availability in YVV
(autotrack): 24.1 kW from 2B & 4B combined.  XPOP (autotrack w/bias):
22.4 kW from both.] 

MCC-H is watching a strong tropical storm coming in from the
Caribbean which may later turn into a hurricane and could require
shifting ISS Flight Control from Houston to Moscow.  A decision
regarding such transfer of control, if required, will in all
likelihood not have to be made before Monday morning (7/14).

 - Continued -

@Message posted automagically by IMTHINGS POST 1.30
--- 
* Origin: SpaceBase(tm) Pt 1 -14.4- Van BC Canada 604-473-9358 (1:153/719.1)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.